OpenStudy (anonymous):

The table shows the number of men and women who have entered a grocery store. What is the experimental probability that the next person to enter the store will be a man?

OpenStudy (anonymous):

Outcome Men Women Frequency 30 25

OpenStudy (goldphenoix):

Are you saying that 25 women and 30 men entered the store?

OpenStudy (anonymous):

yes

OpenStudy (goldphenoix):

Since 25 women entered the store, and 30 men entered the store, you can make a proportion. \[\frac{ 30 }{ 55 }\] 30 represent the men, and the 55 represent the total women and men. This shows you the probability of a man entering a store, 30/55. You can simplify it. \[\frac{ 30 }{ 55 } = \frac{ 6 }{ 11 }\] Answer: The probability of a man entering a store is a 6/11 chance.

OpenStudy (anonymous):

5/6 ?

OpenStudy (goldphenoix):

No, it's 6/11.

OpenStudy (anonymous):

how?

OpenStudy (goldphenoix):

Okay. There's 30 men. There's also 25 women. The denonimator represent the total outcome. 25 is not the total outcome, it's the amount of women that entered the store.

OpenStudy (anonymous):

ohhhhh I got it, 25+30=55 & 30/55= 6/11

OpenStudy (anonymous):

simplify. it.

OpenStudy (goldphenoix):

Think about it. If there was 50 men and there's 5 women. Would you put it as 50/5 or 50/55?. 50/5 would mean there there's a 10/1 chance that there would be a men entering the store. (Which is impossible). 50/55 would simplify to 10/11. Which is reasonable.

OpenStudy (anonymous):

very thanks

OpenStudy (goldphenoix):

I hope this help. :)